OpenStudy (anonymous):

Evaluate the following expression using the values given. Find x + 2y if x = 2 and y = 3

OpenStudy (angela210793):

just put 2 instead of x and 3 instead of y then multiply and add

OpenStudy (angela210793):

try it please....

OpenStudy (anonymous):

would it be 23 or 2 times 3

OpenStudy (angela210793):

it will be 2 times 3

OpenStudy (anonymous):

8 is the answer

OpenStudy (angela210793):

yes :) 2+2*3=2+6=8

OpenStudy (anonymous):

If x=2 and y=3 then x+2y=2+2(3) =2+6 =8